OpenStudy (anonymous):

[Help me - Reward ( Medal and Fan) How much Aluminium will be required to cover a cylinderical flask of length 30 cm??

OpenStudy (anonymous):

the volume should be calculated.. V=pi*R*R*h assuming unit radius V=3.14*1*1*.3=.942

OpenStudy (anonymous):

[2pirh + 2piR^2?\] @MARC_

OpenStudy (anonymous):

cover you mean outer covering? if its the outer covering it should be surface area.. then its 2*pi*r*h + 2*pi*r*r =1.84 + 6.28 =8.122

OpenStudy (anonymous):

The base and lid are not made of aluminium @MARC_ @rohitmathew

OpenStudy (anonymous):

that makes it just the lateral surface.. assuming unit radius.. the answer would be 1.84

OpenStudy (anonymous):

IS THE RADIUS GIVEN?

OpenStudy (butterflydreamer):

oh well if the base and lid aren't made of aluminium it would just be the rectangular side that is aluminium :) So since the length is 30, this means you would use 2pi*r*h so 2pi*r*30 . And then we'd need to know the radius or else we can't solve :)

OpenStudy (anonymous):

The Radius isnt given :( @rohitmathew @butterflydreamer

OpenStudy (butterflydreamer):

what about.. a diagram or the diameter? :O

OpenStudy (anonymous):

u need to put some value for the radius to get a numerical answer.. or the answer would be in terms of the the variable R something like 1.84R..